OpenStudy (anonymous):

I have the answer can someone please check me... The derivative of f(x)=3(cos(2x+4))^7 is f'(x)=ksin(2x+4)(cos(2x+4))^6 For what value of k?

OpenStudy (anonymous):

\[f(x)=3(cos(2x+4))^7 \] \[f'(x)=21\cos(2x+4)^6\sin(2x+4)\times 2=42\cos(2x+4)^6\sin(2x+4)\]

OpenStudy (anonymous):

oops i am off by a minus sign

OpenStudy (anonymous):

\[f'(x)=21\cos(2x+4)^6\times -\sin(2x+4)\times 2=-42\cos(2x+4)^6\sin(2x+4)\]

OpenStudy (anonymous):

So then the value of k is 3

OpenStudy (shamim):

i think u r doing little wrong

OpenStudy (anonymous):

no, \(k=-42\)

OpenStudy (anonymous):

oh ok I understqand I put it into my calculator wrong I just redidi it and I go the right answer now that matches yours

OpenStudy (shamim):

it will b\[f \prime =-3 \times 7 \cos ^{6}(2x+4) \sin (2x+4) \times 2\]

OpenStudy (shamim):

did u c my solution @satellite73